OpenStudy (anonymous):

Given the linear equation y = -4x – 2, find the y-coordinates of the points (1, ), (-1, ), and (-3, ). Please show all of your work

OpenStudy (anonymous):

y = -4x – 2 x=1: y = -4(1) -2 = -6 (1,-6) x=-1: y = -4(-1)-2 = 2 (-1, 2) x=-3: y = -4(-3)-2 = 10 (-3,10)